FLUORESCENCE AND ULTRAVIOLET ABSORPTION SPECTRA AND RING-TWISTING POTENTIAL ENERGY FUNCTIONS OF 1,2-DIHYDRONAPHTHALENE IN ITS S_{0} AND S_{1} STATES

DANIEL E. AUTREY, ZANE A. ARP, and \underline{JAAN LAANE}, Department of Chemistry, Texas A&M University, College Station, TX 77843-3255.